News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 11 February 2005 11.00am EST

During the last week loads were replaned to ensure proper thermal constraints for a propulsion line following a yellow limit violation on 6 Feb. As a result of the replan, the order of targets DLS J0514-4904 and 1249+3038 was changed to ensure a favorable temperature profile.

The schedule of targets for the next week is shown below and includes a DDT observation of PSR J0437-4715.

NGC 3690               ACIS-S       Feb 13
1ES 1239+069           HRC-I        Feb 14
RX J1308.8+2127        ACIS-S
M87                    ACIS-S
Radiation Belts
RX J1308.8+2127        ACIS-S       Feb 15
Cygnus A               ACIS-I
cl1003+3253            ACIS-I
cl1117+1744            ACIS-I       Feb 16
Cygnus A               ACIS-I
Radiation Belts                     Feb 17
RX J0806.4-4123        ACIS-I       Feb 18
PSR J0437-4715         ACIS-S
2MASS 194915-55101     ACIS-S
SDSS J101627.32-00     ACIS-S
RXCJ1213.3-2617        ACIS-S
RX J1308.8+2127        ACIS-S
Cygnus A               ACIS-I       Feb 19
cl0958+4702            ACIS-I
cl0853+5759            ACIS-I
Radiation Belts
HZ43                   HRC-I/LETG   Feb 20
Vega                   ACIS-I
Vega                   ACIS-S
Vega                   HRC-I
Vega (4 obs)           HRC-S


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports